OpenStudy (anonymous):

Two planes, which are 3600 miles part, fly toward each other. Their speeds differ by 50 mph. if they pass each other in 4 hours, what is the speed of each? use the x variable to set up an equation to solve the given problem

OpenStudy (anonymous):

guess we have to use distance equals rate times time

OpenStudy (anonymous):

put one speed as say \(x\) and the faster speed is then \(x+50\) you know their combined speed is therefore \(x+x+50=2x+50\)

OpenStudy (anonymous):

yeah im struggling with this

OpenStudy (anonymous):

then since \(D=RT\) we know \(T=\frac{D}{R}\) and you know \(T=4\), also \(D=3600\) the only thing you don't know is \(x\) solve \[\frac{3600}{2x+50}=4\] for \(x\)

OpenStudy (anonymous):

how would I find the speed of each? @satellite73 do I just sole for x?